OpenStudy (anonymous):

A 50-foot tree casts a shadow 75 feet long. How can you find sine for the angle of elevation that is determined by the sun and the top of the tree?

OpenStudy (anonymous):

\[\sin \left(\tan ^{-1}\left(\frac{50}{75}\right)\right)=\frac{2}{\sqrt{13}}=0.5547 \]